OpenStudy (help!!!!):

the average of 3 numbers is 18. What is 2 times the sum of the 3 numbers?

OpenStudy (anonymous):

so if (x+y+z)/3 = 18 then (x+y+z)= 54 right cause you move the three over by multiplying both sides of the equation by 3 then you can just 54*2 = 108

OpenStudy (anonymous):

make sense?